There are 20 patients in your hospital ward. After an evening meal of BBQ or salad, seven of your patients are quite sick. You suspect something may have been wrong with the BBQ.

  1. Using the table above, what are the odds ratio for the relationship between eating the BBQ and not eating the BBQ? Please show your work.
  2. With your findings, do you believe the BBQ was a possible cause for the sudden sickness in your ward? Explain your reasoning.

See the table below to demonstrate the numbers.

Patients who are sick

Patients who are not sick

Total

Ate the BBQ for dinner

(exposed)

5

a

3

b

8                    

a+b

Did not eat the BBQ for dinner

(unexposed)

2                         

c

10                   

d

12                     

c+d

Total

7                  

a+c

13            

b+d

20          

a+c+b+d
Answer #1

From the details from the above given table we can find out the odds ratio.

Odd ratio= (a*d)/(b*c)

=(5*10)/(3*2)

=50/6

=8.33

8.33 is the odd ratio.

Odd ratio greater than 1 indicate that there is increased probability of occurance of the event.

so I believe BBQ was a possible cause of sudden sickness in the ward.